OpenStudy (anonymous):

Solve the system of linear equations below. Enter your solution as an ordered pair (x,y). Do not include spaces in your answer. 2x+y=7 y=3 I really do suck at linear equations :(

hero (hero):

Hint: y = 3 means that variable y can be replaced with 3 anywhere you see it.

OpenStudy (anonymous):

would it be (2,3)

hero (hero):

In other words, you already have the value of one of the ordered pair coordinates: (x,3) All you need to do now is insert 3 in place of y in 2x + y = 7 to get 2x + 3 = 7 Which if you do so, you get 2x = 7 - 3 x = 4/2 x = 2 So the correct point is (2,3)

OpenStudy (anonymous):

Okay so I was right but thank you I'll remember that

hero (hero):

yw